# cosmICweb: Cosmological Initial Conditions for Zoom-in Simulations in the Cloud

### v1

@@ Abstract
We present the online service cosmICweb (COSMological Initial Conditions on the WEB)— the first database and web interface to store, analyze, and disseminate initial conditions (ICs) for zoom simulations of objects forming in cosmological simulations: from galaxy clusters to galaxies and more. Specifically, we store compressed information about the Lagrangian proto-halo patches for all objects in a typical simulation merger tree along with properties of the halo/galaxy across cosmic time. This enables a convenient web-based selection of the desired zoom region for an object fitting user-specified selection criteria. The information about the region can then be used with the music code to generate the zoom ICs for the simulation. In addition to some other simulations, we currently support all objects in the Eagle simulation database, so that for example the Auriga simulations are easily reproduced, which we demonstrate explicitly. The framework is extensible to include other simulations through an API that can be added to an existing database structure and with which cosmICweb can then be interfaced. We make the web portal and database publicly available to the community.
@@
I Introduction
In cosmology and extragalactic astrophysics, numerical simulations are a fundamental tool to study the growth and evolution of structures in the universe across an enormous range of scales (e.g., Vogelsberger et al., 2020; Angulo & Hahn, 2022, for recent reviews). On the largest scales up to several gigaparsec, they probe the cosmological large scale structure. On the smallest scales ranging from several parsec to kiloparsec, they can resolve the formation and structure of individual galaxies.
The computational cost of computing and storing the full dynamic range in the entire simulated volume sets a limit on the upper and lower scales that can be simulated in a single run. A common approach to circumvent this limitation is the multi-mass or zoom technique, in which the immediate region around the object of interest is simulated with a highly increased resolution compared to its surrounding large scale environment (Katz et al., 1994; Navarro & White, 1994). This method allows capturing both the influences of the cosmic environment from fluctuations on scales equal and larger than the object as well as small scale fluctuations affecting the structure of the object directly. Zoom-in initial conditions can be generated by degrading a high-resolution initial condition outside the zoom region (e.g., Garaldi et al., 2016) or by refining the zoom region (e.g., Bertschinger, 2001; Jenkins, 2010; Hahn & Abel, 2011). Since the former “degrading” approach is computationally limited by memory for deep zoom levels, zoom initial conditions are usually generated using one of the refinement techniques.
When computing such zoom initial conditions, a fundamental ingredient is that the noise field (also known as the random phases) is spatially stable without having to generate it at full resolution. This can be achieved in different ways, with the Panphasia approach (Jenkins, 2013) of an expansion in terms of hierarchically defined basis functions being particularly powerful. Panphasia has been adopted for various large recent simulations such as Eagle (Schaye et al., 2015) and Flamingo (Schaye et al., 2023). Various objects such as e.g., dwarf and Milk-Way-type galaxies in the Auriga simulations (Grand et al., 2024) have been re-simulated at higher resolution using zoom-ins on for example the Eagle volume. However, resimulating objects from these simulations so far required proprietary software and access to the particle data at initial and final time. Often, the parent simulations of zoom simulations are not even published in other studies.
@@ I Introduction
The steps 1 to 5 are preparatory steps that usually have to be repeated for new zoom simulation, unless they are based on the same parent volume. Public halo and galaxy catalogs of large simulations exist (e.g., McAlpine et al., 2016; Nelson et al., 2019; Heitmann et al., 2019), but they are usually not associated with proto-halo patches. Furthermore, the absence of public access to initial conditions in research articles complicates the direct comparison between different simulation codes, baryonic physics prescriptions, and the verification of scientific results.
@@
  • Referencing initial conditions. Users can create references to halos and collections of halos that they used in their works, allowing the community to easily reproduce results and to compare different simulation codes for verification. For each such reference, a tag is generated that can be included in research articles, allowing the community to access the zoom initial conditions using the cosmICweb application or directly through a plugin distributed with music.
@@
This paper accompanies the release of cosmICweb 222cosmICweb URL: https://cosmicweb.eu. and documents the simulations and pre-computed initial conditions that are available at release and the architecture and extendability of cosmICweb. The structure is as follows: In Section II, we introduce the simulations that are included in the cosmICweb application at release and give an overview of the available halo, galaxy and Lagrangian ellipsoid selectors which can be used to search the simulation databases. In Section III, we discuss the computation of proto-halo patches from simulation data, in particular our choice of minimum bounding ellipsoids to describe Lagrangian proto-halos. In Section IV, we present the individual components of cosmICweb and how its interface can be used to find, download, and publish initial conditions for zoom simulations. In Section V, we analyze the proto-halo patches provided on cosmICweb by measuring their size and efficiency distributions depending on halo and environmental properties and the choice of traceback-radius. We also show two proof-of-concept zoom simulations to highlight the capabilities of cosmICweb. We conclude in Section VI. More details on the requirements of connecting additional simulations to cosmICweb can be found in Appendix A.
II Cosmological simulations and merger trees
@@
The cosmICweb platform has been designed in a modular way to allow for the integration of data from diverse projects hosted at different locations (also see Section IV.1). Therefore, more simulations can easily be added in the future by us and the community. Here, we will briefly summarize the simulations that are available at the date of release.
@@
Table 1: List of simulations available at the time of release of cosmICweb. The simulation parameters include the box length of the simulation volume, the cosmological parameters that were used, dark matter mass resolution, number of snapshots, the structure finder that was used, and the minimum number of particles within the traceback-radius that we require to compute the minimum bounding ellipsoids, $N^{e}_{\min}$.
box length cosmology DM resolution galaxies $N^{e}_{\min}$ # snapshots halo finder
$h^{-1}\mathrm{Mpc}$ ] $h^{-1}\mathrm{M}_{\odot}$ ] [redshift range]
primary agora 60 Planck 2015 $1.21\times 10^{8}$ 1000 101 [12.0 – 0] rockstar
150MPC 150 Planck 2015 $2.70\times 10^{8}$ 100 101 [12.0 – 0] rockstar
300MPC 300 Planck 2015 $2.14\times 10^{9}$ 100 101 [12.0 – 0] rockstar
Rhapsody 1000 WMAP 7/9 $6.46\times 10^{10}$ 1000 101 [12.0 – 0] rockstar
Rhapsody_PlanckCosmo 1000 Planck 2015 $7.99\times 10^{10}$ 1000 101 [12.0 – 0] rockstar
EAGLE Ref-L0100N1504 $100h$ Planck 2014 $9.7\times 10^{6}h$ 1000 29 [20.3 – 0] fof/subfind
Ref-L0025N0376 $25h$ Planck 2014 $9.7\times 10^{6}h$ 1000 29 [20.3 – 0] fof/subfind
Ref-L0025N0752 $25h$ Planck 2014 $1.21\times 10^{6}h$ 1000 29 [20.3 – 0] fof/subfind
DMONLY-L0100N1504 $100h$ Planck 2014 $1.15\times 10^{7}h$ 1000 29 [20.3 – 0] fof/subfind
DMONLY-L0025N0376 $25h$ Planck 2014 $1.15\times 10^{7}h$ 1000 29 [20.3 – 0] fof/subfind
DMONLY-L0025N0752 $25h$ Planck 2014 $1.44\times 10^{6}h$ 1000 29 [20.3 – 0] fof/subfind
II.1 Gravity-only simulations
@@
Currently, we are locally hosting data from five gravity-only simulations. The parameters that were used for these simulations are summarized in Table 1. The simulations span cosmological volumes from $60^{3}$ $h^{-3}$Mpc3 to 11 $h^{-3}$Gpc3 and use cosmological parameters conforming with the Planck 2015 (Planck Collaboration, 2015) and WMAP 7/9 (Komatsu et al., 2011; Hinshaw et al., 2013) results. We include two simulations which have been previously used for zoom simulations: the cosmological volumes of the AGORA comparison project (Kim et al., 2013) and of the RHAPSODY cluster re-simulation project (Wu et al., 2013b, a; Hahn et al., 2017) which is originally based on the Carmen simulation of the LasDamas project (McBride et al., 2009) – we labeled the simulations agora and Rhapsody for the sake of convenience. Rhapsody uses older WMAP 7/9 cosmological parameters. However, we have updated the parameters in the Rhapsody_PlanckCosmo simulation to allow comparisons between the halo catalogs and zoom simulation results.
@@
Initial conditions for the full simulations were generated using music (Hahn & Abel, 2011) at redshift $z=49$ (Rhapsody & Rhapsody_PlanckCosmo) and $z=99$ (remaining simulations). Using a modified version of the gadget-3 code (Springel, 2005), we evolved the simulations to redshift $z=0$ and stored 100 snapshots between $z=12$ and 0, logarithmically distributed in scale factor units. Each snapshot was processed with the rockstar structure finder (Behroozi et al., 2013a), which identifies halos and subhalos using a six-dimensional phase-space friend-of-friend algorithm and measures their properties, including spherical overdensity masses, radii, and concentrations, spin parameters, and relaxedness quantities. Using consistent-trees (Behroozi et al., 2013b), we group halos across snapshots to merger trees containing information about spatial hierarchy (halo substructure) as well as temporal hierarchy (descendant and progenitor halos and merger events). In addition, we compute environmental parameters such as the normalized distance to the closest more massive neighbor $D_{1,1}$ (cf. Table 2) directly from the halo catalog.
@@
For every host halo in each snapshot, we compute the Lagrangian minimum bounding ellipsoid at $1,2,4$, and $10$ $R_{\mathrm{vir}}$, if the number of particles within that radius exceeds the threshold listed in Table 1. We will discuss the definition, construction, and statistical distribution of these proto-halo patches in more detail in Section III and Section V.1.
@@
The halo catalogs including the merger-tree information and the Lagrangian proto-halos are stored in a relational database. To reduce latency and speed-up queries, the data has been partitioned and indexed on the most commonly used columns. More details on the storage of and interface to the data can be found in Section IV.3.
II.2 The Eagle database
@@
The phases for the initial conditions for the Eagle simulations were extracted from regions within the large Panphasia multi-scale white noise field (Jenkins, 2013). The phases for the Eagle volumes are published in Schaye et al. (2015). The updated version of music  [^4] The updated version 2 of music is available from
@@
includes the Panphasia white noise field, allowing us to create zoom simulations in the Eagle volumes from a unified interface.
@@
The host halos for the Auriga Project galaxy formation simulations (Grand et al., 2017, 2024) were selected from the Eagle 100 Mpc side-length box. Table A1 in Grand et al. (2017) matches the host halos of the Auriga sample with their corresponding halo IDs in the Eagle database. In Section V.3, we recreate the host halo of the Auriga 6 galaxy and compare the redshift zero dark matter halo properties against an original Auriga project dark matter simulation of that same halo.
III Proto-halo patches
@@
Figure 2: Slices through six halos at redshift $z=0$ (right panels) and their proto-halo patches (left panels). The halos are selected from the 300MPC simulation, centered around three halo mass values $M_{200c}\sim 10^{12},10^{13}$, and $10^{14}$ $h^{-1}\mathrm{M}_{\odot}$ from isolated environments (top) and more populated regions (bottom). The circles in Eulerian space correspond to traceback-radii of 1 (magenta), 2 (green), 4 (blue), and 10 $R_{\mathrm{vir}}$ (yellow), respectively. Particles within $R_{\mathrm{tb}}$ are traced back to Lagrangian space and colored correspondingly. The ellipses represent slices through the minimum bounding ellipsoids and the background shading illustrate the density field (darker colors correspond to higher density).
@@
In this section, we present an overview of the zoom regions available in cosmICweb. Specifically, we define Lagrangian proto-halos and explain the rationale behind our choice of minimum bounding ellipsoids to describe these volumes.
@@
The proto-halo patch is defined by the Lagrangian volume that collapses to a halo as the universe evolves. In an $N$-body simulation, it is traced by the particles that are within the traceback-radius $R_{\mathrm{tb}}$ of the halo, which is often chosen as a multiple of the halo radius defined by a spherical overdensity criterion. Choosing a larger traceback-radius increases the computational cost of the simulation (although not proportionally to the volume, as the density decreases with the distance to the halo), but reduces the probability of contamination by low-resolution particles during the evolution of the halo. A larger zoom region also allows to resolve the immediate environment of the halo in high resolution, which is desirable in many simulations. However, a too large traceback radius potentially includes surrounding halos, in particular if the targeted halo is located in a populated environment, resulting in a computationally more expensive simulation.
@@
Figure 2 illustrates the relation between proto-halos in Lagrangian space and the resulting halo at redshift $z=0$. We choose six halos from the 300MPC simulation, two halos each around $M_{200c}\sim 10^{12},10^{13}$, and $10^{14}$ $h^{-1}\mathrm{M}_{\odot}$, with one from an isolated and one from a more populated environment. The traceback radii at 1, 2, 4, and 10 $R_{\mathrm{vir}}$ are shown as circles in Eulerian space. All particles within the circles are traced back to the initial conditions and colored with the color corresponding to $R_{\mathrm{tb}}$. We note that while in isolated environments, the proto-halo patches are compact and do not grow significantly with increasing traceback radius (with the exception of $R_{\mathrm{tb}}=10R_{\mathrm{vir}}$), they are less spherical, more scattered and grow significantly at larger radii in populated regions.
@@
In general, the efficiency increases with the complexity of the boundary description, e.g the rectangular bounding box requiring 6 coordinates is less efficient than the minimum bounding ellipsoid (9 parameters), which in turn is less efficient than the convex hull described by all the particles on the surface (Oñorbe et al., 2013). As we will show in Section V.1, the variance of the efficiency is mass dependent and increases for low-mass halos due to environmental effects, as most of the low-mass halos evolve in clustered environments with strong tidal fields. In contrast, more massive halos form at the nodes of the cosmic web where mass accretion occurs more isotropically, leading to a more well-behaved, spherical proto-halo.
The efficiency parameter is a measure of how well the abstracted Lagrangian volume describes the proto-halo patch and determines the fraction of computational resources that is focused on the volume of interest within the traceback-radius. It is therefore desirable to create proto-halo patches with high $\mathcal{E}$ for zoom simulations. The efficiency parameter naturally depends on the environment and formation history of the halo, which is why one has to be careful to not create a biased sample when selecting halos by the efficiency parameters of their Lagrangian minimum bounding ellipsoid. We will discuss this bias more quantitatively in Section V.2.
@@
III.1 Minimal bounding ellipsoids in cosmICweb
For cosmICweb, we provide proto-halo regions in form of minimal bounding ellipsoids. This choice is motivated by the impracticality of storing convex hulls for all halos across all snapshots and multiple traceback-radii. Minimal bounding ellipsoids offer a robust alternative while still capturing the overall shape of typical proto-halo patches at manageable data storage requirements.
An ellipsoid is described by its center $\bm{q}_{c}$ and its $3\times 3$ shape matrix $\mathrm{A}$. The symmetric, positive definite matrix $\mathrm{A}$ can be normalized such that coordinates $\bm{q}$ that are within the ellipsoid satisfy

### v3

@@ Abstract
We present the online service cosmICweb (COSMological Initial Conditions on the WEB)— the first database and web interface to store, analyze, and disseminate initial conditions (ICs) for zoom simulations of objects forming in cosmological simulations: from galaxy clusters to galaxies and more. Specifically, we store compressed information about the Lagrangian proto-halo patches for all objects in a typical simulation merger tree along with properties of the halo/galaxy across cosmic time. This enables a convenient web-based selection of the desired zoom region for an object fitting user-specified selection criteria. The information about the region can then be used with the music code to generate the zoom ICs for the simulation. In addition to some other simulations, we currently support all objects in the Eagle simulation database, so that for example the Auriga simulations are easily reproduced, which we demonstrate explicitly. The framework is easily extensible to include other simulations through an API that can be added to an existing database structure and with which cosmICweb can then be interfaced. We make the web portal and database publicly available to the community.
thanks: E-mail: mbuehlmann@anl.gov
1 Introduction
In cosmology and extragalactic astrophysics, numerical simulations are a fundamental tool to study the growth and evolution of structures in the universe across an enormous range of scales (e.g., Vogelsberger et al., 2020; Angulo & Hahn, 2022, for recent reviews). On the largest scales up to several gigaparsec, they probe the cosmological large scale structure. On the smallest scales ranging from several parsec to kiloparsec, they can resolve the formation and structure of individual galaxies. The computational cost of computing and storing the full dynamic range in the entire simulated volume sets a limit on the upper and lower scales that can be simulated in a single run. A common approach to circumvent this limitation is the multi-mass or zoom technique, in which the immediate region around the object of interest is simulated with a highly increased resolution compared to its surrounding large scale environment (Katz et al., 1994; Navarro & White, 1994). This method allows capturing both the influences of the cosmic environment from fluctuations on scales equal and larger than the object as well as small scale fluctuations affecting the structure of the object directly. Zoom-in initial conditions can be generated by degrading a high-resolution initial condition outside the zoom region (e.g., Garaldi et al., 2016) or by refining the zoom region (e.g., Bertschinger, 2001; Jenkins, 2010; Hahn & Abel, 2011). Since the former “degrading” approach is computationally limited by memory for deep zoom levels, zoom initial conditions are usually generated using one of the refinement techniques.
@@ 1 Introduction
The steps 1 to 5 are preparatory steps that usually have to be repeated for new zoom simulation, unless they are based on the same parent volume. Public halo and galaxy catalogs of large simulations exist (e.g., McAlpine et al., 2016; Klypin et al., 2017; Nelson et al., 2019; Heitmann et al., 2019), but they are usually not associated with proto-halo patches. Furthermore, the absence of public access to initial conditions in research articles complicates the direct comparison between different simulation codes, baryonic physics prescriptions, and the verification of scientific results.
@@
  • Referencing initial conditions. Users can create references to halos and collections of halos that they used in their works, allowing the community to easily reproduce results and to compare different simulation codes and physics implementations. For each such reference, a tag is generated that can be included in research articles, allowing the community to access the zoom initial conditions using the cosmICweb application or directly through a plugin distributed with music.
@@
This paper accompanies the release of cosmICweb 222cosmICweb URL: https://cosmicweb.eu. and documents the simulations and pre-computed initial conditions that are available at release and the architecture and extendability of cosmICweb. The structure is as follows: In Section 2, we introduce the simulations that are included in the cosmICweb application at release and give an overview of the available halo, galaxy and Lagrangian ellipsoid selectors which can be used to search the simulation databases. In Section 3, we discuss the computation of proto-halo patches from simulation data, in particular our choice of minimum bounding ellipsoids to describe Lagrangian proto-halos. We also present an analysis of the proto-halos provided on cosmICweb by measuring their size and efficiency distributions depending on halo and environmental properties and the choice of traceback-radius. In Section 4, we present the individual components of cosmICweb and how its interface can be used to find, download, and publish initial conditions for zoom simulations. In Section 5, we show two proof-of-concept zoom simulations to highlight the capabilities of cosmICweb. We conclude in Section 6. More details on the requirements of connecting additional simulations to cosmICweb can be found in Appendix B.
2 Cosmological simulations and merger trees
@@
The cosmICweb platform has been designed in a modular way to allow for the integration of data from diverse projects hosted at different locations (also see Section 4.1). Therefore, more simulations can easily be added in the future by us and the community. Here, we will briefly summarize the simulations that are available at the date of release.
@@
Table 1: List of simulations available at the time of release of cosmICweb. The simulation parameters include the box length $L$ of the simulation volume, the cosmological parameters that were used, dark matter mass resolution $m_{\mathrm{DM}}$, gravitational force resolution $\epsilon$, number of snapshots, the structure finder that was used, and the minimum number of particles within the traceback-radius that we require to compute the minimum bounding ellipsoids, $N^{e}_{\min}$.
$L$ cosmology $m_{\mathrm{DM}}$ $\epsilon_{\mathrm{com}}$ $N^{e}_{\min}$ snapshots halo finder G
$h^{-1}\mathrm{Mpc}$ $h^{-1}\mathrm{M}_{\odot}$ $h^{-1}\mathrm{kpc}$ [redshift range]
primary agora 60 Planck 2015 $1.21\times 10^{8}$ 4 1000 101 [12.0 – 0] rockstar
150MPC 150 Planck 2015 $2.70\times 10^{8}$ 5 100 101 [12.0 – 0] rockstar
300MPC 300 Planck 2015 $2.14\times 10^{9}$ 10 100 101 [12.0 – 0] rockstar
Rhapsody 1000 WMAP 7/9 $6.46\times 10^{10}$ 50 1000 101 [12.0 – 0] rockstar
Rhapsody_PlanckCosmo 1000 Planck 2015 $7.99\times 10^{10}$ 50 1000 101 [12.0 – 0] rockstar
[…]
  • $\dagger$
    This column indicates simulations that have been run with baryonic physics.
2.1 Gravity-only simulations
@@
Currently, we are locally hosting data from five gravity-only simulations. The parameters that were used for these simulations are summarized in Table 1. The simulations span cosmological volumes from $60^{3}$ $h^{-3}$pc3 to $1$ $h^{-3}$pc3 and use cosmological parameters conforming with the Planck 2015 (Planck Collaboration, 2015) and WMAP 7/9 (Komatsu et al., 2011; Hinshaw et al., 2013) results. We include two simulations which have been previously used for zoom simulations: the cosmological volumes of the AGORA comparison project (Kim et al., 2013) and of the RHAPSODY cluster re-simulation project (Wu et al., 2013b, a; Hahn et al., 2017) which is originally based on the Carmen simulation of the LasDamas project (McBride et al., 2009) – we labeled the simulations agora and Rhapsody for the sake of convenience. Rhapsody uses older WMAP 7/9 cosmological parameters. However, we have updated the parameters in the Rhapsody_PlanckCosmo simulation to allow comparisons between the halo catalogs and zoom simulation results.
Initial conditions for the full simulations were generated using music (Hahn & Abel, 2011) at redshift $z=49$ (Rhapsody & Rhapsody_PlanckCosmo) and $z=99$ (remaining simulations). Using the L-gadget-3 code – a lean version of the gadget-3 code developed for the Millennium-XXL simulation (Angulo et al., 2012) – we evolved the simulations to redshift $z=0$ and stored 100 snapshots between $z=12$ and 0, logarithmically distributed in scale factor units. We adopted a fixed co-moving softening of $1/30$th of the linear particle separation for the “primary” runs. The numerical values of the force softening lengths for the “primary” and EAGLE simulations are listed in Table 1.
@@
Each snapshot was processed with the rockstar structure finder (Behroozi et al., 2013a), which identifies halos and subhalos using a six-dimensional phase-space friend-of-friend algorithm and measures their properties, including spherical overdensity masses, radii, and concentrations, spin parameters, and relaxedness quantities. Using consistent-trees (Behroozi et al., 2013b), we group halos across snapshots to merger trees containing information about spatial hierarchy (halo substructure) as well as temporal hierarchy (descendant and progenitor halos and merger events). In addition, we compute environmental parameters such as the normalized distance to the closest more massive neighbor $D_{1,1}$ (cf. Table 3) directly from the halo catalog.
@@
For every host halo in each snapshot, we compute the Lagrangian minimum bounding ellipsoid at $1,2,4$, and $10$ $R_{\mathrm{vir}}$, if the number of particles within that radius exceeds the threshold listed in Table 1. We will discuss the definition, construction, and statistical distribution of these proto-halo patches in more detail in Section 3 and Section 3.2.
@@
The halo catalogs including the merger-tree information and the Lagrangian proto-halos are stored in a relational database. To reduce latency and speed-up queries, the data has been partitioned and indexed on the most commonly used columns. More details on the storage of and interface to the data can be found in Section 4.3.
2.2 The Eagle database
@@
[…]
@@
[…]
@@
The host halos for the Auriga Project galaxy formation simulations (Grand et al., 2017, 2024) were selected from the Eagle 100 Mpc side-length box. Table A1 in Grand et al. (2017) matches the host halos of the Auriga sample with their corresponding halo IDs in the Eagle database. In Section 5, we recreate the host halo of the Auriga 6 galaxy and compare the redshift zero dark matter halo properties against an original Auriga project dark matter simulation of that same halo.
3 Proto-halo patches
@@
Figure 2: Slices through six halos at redshift $z=0$ (right panels) and their proto-halo patches (left panels). The halos are selected from the 300MPC simulation, centered around three halo mass values $\smash{M_{200,c}\sim 10^{12},10^{13}}$, and $\smash{10^{14}}$ $h^{-1}\mathrm{M}_{\odot}$ from isolated environments (top) and more populated regions (bottom). The circles in Eulerian space correspond to traceback-radii of 1 (magenta), 2 (green), 4 (blue), and 10 $R_{\mathrm{vir}}$ (yellow), respectively. Particles within $R_{\mathrm{tb}}$ are traced back to Lagrangian space and colored correspondingly. The ellipses represent slices through the minimum bounding ellipsoids and the background shading illustrate the density field (darker colors correspond to higher density).
@@
In this section, we present an overview of the zoom regions available in cosmICweb. Specifically, we define Lagrangian proto-halos and explain the rationale behind our choice of minimum bounding ellipsoids to describe these volumes. We also present an analysis of the ellipsoids’ mass and environment dependent efficiency, i.e. how well the ellipsoids are capturing the proto-halo shape. We measure the mass growth of the ellipsoids depending on the size of the zoom region and quantify the selection bias that is introduced if constraints on the efficiency are imposed.
Generally, a cosmological $N$-body simulation maps a fluid element, or $N$-body particle, from its initial unperturbed ‘Lagrangian’ position $\bm{q}$ along a trajectory to a position $\bm{x}(\bm{q},z)$ in the evolved universe at redshift $z$. By definition thus, the Lagrangian coordinate $\bm{q}$ serves as a unique label of each particle over the course of a simulation. Tracing back the set of all particles $\mathcal{P}(\mathcal{H})=\{\bm{q}_{i}\}_{\bm{x}(\bm{q},z)\in\mathcal{H}}$ making up a halo $\mathcal{H}$ yields the proto-halo patch $\mathcal{P}$, i.e. the region of primordial Universe that will collapse to form this halo. In general, the set $\mathcal{P}$ will have irregular shape, and potentially even holes.
Since the definition of a halo is to some degree arbitrary (see e.g. the discussion in White (2001)) and, depending on the application, one in many cases is not only interested in a specific halo but also its surroundings, we allow for a range of traceback-radii centered at halo locations. This traceback-radius $R_{\mathrm{tb}}$ of the halo, is chosen as a multiple of the halo radius defined by a spherical overdensity criterion.
@@
Choosing a larger traceback-radius increases the computational cost of the simulation (although not proportionally to the volume, as the density decreases with the distance to the halo), but reduces the probability of contamination by low-resolution particles during the evolution of the halo. A larger zoom region also allows to resolve the immediate environment of the halo in high resolution, which is desirable in many simulations. However, a too large traceback radius potentially includes surrounding halos, in particular if the targeted halo is located in a populated environment, resulting in a computationally more expensive simulation.
@@
Figure 2 illustrates the relation between proto-halos in Lagrangian space and the resulting halo at redshift $z=0$. We choose six halos from the 300MPC simulation, two halos each around $M_{200,c}\sim 10^{12},10^{13}$, and $10^{14}$ $h^{-1}\mathrm{M}_{\odot}$, with one from an isolated and one from a more populated environment. The traceback radii at 1, 2, 4, and 10 $R_{\mathrm{vir}}$ are shown as circles in Eulerian space. All particles within the circles are traced back to the initial conditions and colored with the color corresponding to $R_{\mathrm{tb}}$. We note that while in isolated environments, the proto-halo patches are compact and do not grow significantly with increasing traceback radius (with the exception of $R_{\mathrm{tb}}=10R_{\mathrm{vir}}$), they are less spherical, more scattered and grow significantly at larger radii in populated regions.
@@
In general, the efficiency increases with the complexity of the boundary description, e.g the rectangular bounding box requiring 6 coordinates is less efficient than the minimum bounding ellipsoid (9 parameters), which in turn is less efficient than the convex hull described by all the particles on the surface (Oñorbe et al., 2013). As we will show in Section 3.2, the variance of the efficiency is mass dependent and increases for low-mass halos due to environmental effects, as most of the low-mass halos evolve in clustered environments with strong tidal fields. In contrast, more massive halos form at the nodes of the cosmic web where mass accretion occurs more isotropically, leading to a more well-behaved, spherical proto-halo.
The efficiency parameter is a measure of how well the abstracted Lagrangian volume describes the proto-halo patch and determines the fraction of computational resources that is focused on the volume of interest within the traceback-radius. It is therefore desirable to create proto-halo patches with high $\mathcal{E}$ for zoom simulations. The efficiency parameter naturally depends on the environment and formation history of the halo, which is why one has to be careful to not create a biased sample when selecting halos by the efficiency parameters of their Lagrangian minimum bounding ellipsoid. We will discuss this bias more quantitatively in Section 3.3.
@@
3.1 Minimal bounding ellipsoids in cosmICweb
For cosmICweb, we provide proto-halo regions in form of minimal bounding ellipsoids. This choice is motivated by the impracticality of storing convex hulls for all halos across all snapshots and multiple traceback-radii. Minimal bounding ellipsoids offer a robust alternative while still capturing the overall shape of typical proto-halo patches at manageable data storage requirements. They furthermore provide a conservative choice in the sense that they are guaranteed to fill in potential holes in the Lagrangian patch.
An ellipsoid is described by its center $\bm{q}_{c}$ and its $3\times 3$ shape matrix $\mathrm{A}$. The symmetric, positive definite matrix $\mathrm{A}$ can be normalized such that coordinates $\bm{q}$ that are within the ellipsoid [^5] A matrix is positive definite if one can write $\mathrm{A}=\mathrm{B}^{\top}\mathrm{B}$. The ellipsoid equation can thus also be written as that for a ball $\|\bm{p}\|^{2}<1$ where $\bm{p}:=\mathrm{B}(\bm{q}-\bm{q}_{c})$ represents the inverse deformation of the ellipsoid to the sphere so that the eigenvalues of $\mathrm{B}$ are just the inverse principal axis lengths. satisfy